OpenStudy (anonymous):

x2 - 1 = x + 1 (1 point) x = -1 x = -2 x = 2 and x = -1 x = -1 and x = -2

OpenStudy (anonymous):

420 420

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

x^2-1=x+1 x^2-1-x-1=0 x^2-x-2=0 (x-2)(x+1)=0 x=2 or x=-1

OpenStudy (anonymous):

third one :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcomev :)